dn 6. Use the theorem, L {t f(t)} = (-1) den! F(s), to find each of the following: (a) L {t cos 2t} (b) L {t sint}
Answer #1

6 6) a) L{ t cus2t} 1 L { ť cos (26)] (-) d ds [L{cos (2+)} d ds ( s²+2² (sa) () sd (sta) (544)2 - 8²4 4 - 25² (574) 2 2 S2. 3 (4) م (5 4 } (ط (1) de [L{sines 2. d² 1 2. ds? | } کی ارچ -2 d - بم 2 2 ds d ds هما } الم و 2 = ds ( 51) 2 2. (۱) d as dدارد ؟ - (و) ی را پی) 2- = S P sp SP (۱۱۹ 2. (25) (61) s02 - 2 (۱ی - 2x (5) 2 ہ 2. - X 7 - و 4 - 1 و 2 (1) 4 = (ری) 24 ار3 -